OpenStudy (anonymous):

Need help figuring out how to do this Please. Indicate in standard form the equation of the line passing through the given points. B(0, 0), C(4, -4)

OpenStudy (jhannybean):

Use point slope form \[y-y _{1}=m(x-x _{1})\]

OpenStudy (anonymous):

y - y1 = m(x - x1) m is slope m = (y2-y1)/(x2-x1)

OpenStudy (jhannybean):

to figure out m, use the two points B and C, in the formula: \[\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\]

OpenStudy (anonymous):

ok I think I got it

OpenStudy (jhannybean):

Good :) glad to help

OpenStudy (anonymous):

thanks much :)

OpenStudy (jhannybean):

Yw